OpenStudy (anonymous):

the probability that a professional baseball player will get a hit is 1/3. calculate the exact probability that he will get at least 3 hits in 5 attempts

OpenStudy (anonymous):

is it 17/81 ??

OpenStudy (anonymous):

I got that too however, it's actually 51/243 i think the 17/1 is considered as 1 hit, so multiply that by 3. THat's what the regents book gave me

OpenStudy (anonymous):

aren't the two fractions same :P

OpenStudy (anonymous):

You have to use binomial distribution here, success: p=(1/3) failure: q=1-p=(2/3) So for at least 3 hits its 10*(1/3)^3*(2/3)^2=40/243=0.164 so over 16%

OpenStudy (anonymous):

they are but honestly i don't think they'd want you to simplify :O then again i'm not so sure!

OpenStudy (anonymous):

@Arthur_Bonnouvrier - yeah but you will have to add all the probabilities right... P(3)+P(4)+P(5)... the number of hits are in bracket...

OpenStudy (anonymous):

@mistuh - i dont think it matters if you simplify or not... they are just fractions... neways if your book says not to simplify then ok... Using the binomial method you get 51/243 not 17/81... But i still think it doesn't matter...

OpenStudy (anonymous):

I'm brushing up before my regents exam today, don't know much about anything :P